How many four digit numbers can be configured using 0-9 if the number must be at least 5000 and divisible by 10?
Answer by devipriya(2)

The first digit must be 5, 6, 7, 8 or 9.
There are 9P3 permutations of the remaining numbers taken 3 at a time. Combined with the 5 numbers allowed for the most significant digit this gives 5*9P3 configurations.
However the least significant digit must be 0 to enable division by 10, disallowing 9 numbers.
Therefore the allowable number of configurations is:
=(5*9p3)/9=(5*9*8*7)/9
=280